What is the solution of the square root of two x plus four (square root bracket thing over two x plus four) equals x minus two?
Answer by unlockmath(1614)

Hello,
Let's write this out as:
sq rt(2x+4)=x-2
Square both sides:
2x+4=(x-2)^2
Expand this out:
2x+4=x^2-4x+4
Subtract 2x and 4:
0=x^2-6x
Factor:
x(x-6)
Solve for x:
x=0
x=6
check each answer to see if they work:
0 doesn't work
6 does.
Make sense?
RJ
www.math-unlock.com